OpenStudy (beccab003):

Can someone please check my answer? It is an easy question. I will medal the person with the best answer. Thanks! You want to get rid of the y by elimination in the system below. x – 2y = -1 2x – 3y = 4 If you multiply the top equation by 6, by which number will you multiply the bottom equation? 4 or -3 -4 3 or -3 -4 or 3 I think it is -4.

OpenStudy (beccab003):

@Hero

hero (hero):

What what you get if you multiplied the top equation by 6?

hero (hero):

Me personally, I wouldn't multiply the top equation by 6 to eliminate this, but hey, that's what the problem calls for.

OpenStudy (beccab003):

6x - 12y = -6

OpenStudy (beccab003):

I have already done the problem and I know how to do it. I was just checking to make sure I did it right.

hero (hero):

Yes, you are correct. You multiply the second equation by -4 to eliminate y

OpenStudy (beccab003):

Okay thanks! Would you mind checking one more problem?

hero (hero):

As long as it's quick.

OpenStudy (beccab003):

okay. here:What does x equal in the solution of the system of equations below? X + 2y + z = 4 4y – 3z = 1 2y + 10z = 12 7 3 5 1

hero (hero):

Divide both sides of the third equation by 2. What do you get?

OpenStudy (beccab003):

1/2y + 5z = 6

OpenStudy (beccab003):

sorry

OpenStudy (beccab003):

y = -5z +6

hero (hero):

Substitute the expression for y into the second equation of your system

OpenStudy (beccab003):

So I found out that z = 1

OpenStudy (beccab003):

and y =

OpenStudy (beccab003):

1 too

hero (hero):

Whatever you found, show the work you did to find it.

OpenStudy (beccab003):

4(-5z+6)-3z=1

hero (hero):

Show the steps as well

OpenStudy (beccab003):

-20z+24-3z=1

OpenStudy (beccab003):

I am showing you the steps now

hero (hero):

Just type all the steps all at once.

OpenStudy (beccab003):

4(-5z+6)=-3z=1 -20z+24-3z=1 -23z+24=1 z=1

OpenStudy (beccab003):

y=-5(1)+6

OpenStudy (beccab003):

y=1

hero (hero):

Okay, that looks good. If you found y and z, then you should be able to find the value of x easily.

OpenStudy (beccab003):

x+2(1)+1=4 x+3=4 x=1

OpenStudy (beccab003):

I that correct?

hero (hero):

Looks right

OpenStudy (beccab003):

Thanks for you help!

OpenStudy (anonymous):

looks good to me